OpenStudy (anonymous):

Find all real number solutions 3x^-2+7x^-1+2=0 ...

OpenStudy (anonymous):

what if i find artificial solutions instead of real :P

OpenStudy (anonymous):

lol i dont think that will work... but i sure wish!

OpenStudy (anonymous):

multiply by x^2 both sides

OpenStudy (anonymous):

you'll get: 2x^2+7x+3=0 solve it

OpenStudy (anonymous):

x=-3 & x=-1/2

OpenStudy (anonymous):

k thanks! ill try and see if i can get the same!

OpenStudy (anonymous):

Good luck:)

myininaya (myininaya):

u=x^(-1) u^2=x^(-1*2)=x^(-2) 3u^2+7u+2=0 3u^2+6u+u+2=0 3u(u+2)+(u+2)=0 (u+2)(3u+1)=0 u=-2 u=-1/3 but u=x^(-1) x^(-1)=-2 x=-1/2 x^(-1)=-1/3 x=-3

OpenStudy (anonymous):

THANK YOU

OpenStudy (anonymous):

a quadratic in \[ \frac{1}{x}\]

OpenStudy (anonymous):

oh myininaya got it! no surpise

myininaya (myininaya):

lol

OpenStudy (anonymous):

now if would only write in latex i could read it!

myininaya (myininaya):

satellite i want to show off i want to show you my logic skills follow me if you can

OpenStudy (anonymous):

are you posting a question?

myininaya (myininaya):

no you have to go to my name on the homepage and quick where im at hope it works such a lag site

OpenStudy (anonymous):

i;ll try